OpenStudy (anonymous):

Why is the formulae for force as a result of electromagnetic forces given by:

OpenStudy (anonymous):

\[F=q(E+\frac{v}{c} \times B)\]

OpenStudy (kociela):

first one comes how we have measured the electric field; so E = F / q what in turns leads to coulomb law ( not sure which one was first, but the coulomb law is derived from experimental results, theoretical should exist but don't know is there one yet ) In the second one i think it should be F = q * v x B and and can be formulated to E = v x B this gives the E and B perpendicular by each other which holds experimentally. that looks like it could be derived from equation compilation by Maxwell and trigonometric wave forms ( A_c = A * sin k(x-v*t) ) which would give experimental background in the end. Forward from this to theoretical backgrounds goes little over my head by now. But it's possible.